# What is Fair? Defining Fairness in Machine Learning for Health

### v3

@@ Abstract
In predicting hospital readmission, the probability of being readmitted given the model makes that prediction and the probability of not being readmitted given the model makes that prediction is the same for Black and white patients [94].
Predictive Parity $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospital readmission, the probability of being readmitted given the model makes that prediction is the same for Black and white patients [94].
Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all values of the score for males and females [95].
Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all values of the score [95].
@@
Brier Score Parity $\mathbb{E}[(Y-S)^{2}\mid A=a_{0}]=\mathbb{E}[(Y-S)^{2}\mid A=a_{1}],\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to enroll patients into high-risk care management programs, the mean squared error between the score and the label is the same for Black and White patients [16].
Overall Accuracy Equality $P(D=Y\mid A=a_{0})=P(D=Y\mid A=a_{1})\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the probability of a correct classification (i.e., correctly classifying a patient as having or not having lung cancer) is the same for males and females [93].
Treatment Equality #$\frac{\#\{D=0,Y=1,A=a_{0}\}}{\#\{D=1,Y=0,A=a_{0}\}}=\frac{\#\{D=0,Y=1,A=a_{1}\}}{\#\{D=1,Y=0,A=a_{1}\}}\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the ratio of the false negatives and false positives is the same for both males and females [93].
Table 3: Common group fairness criteria. Mathematical definitions of group fairness criteria, their interpretation in the context of an example from the literature, and an indication of whether the criterion is a relaxation of independence ($D\perp\!\!\!\perp A$), separation ($D\perp\!\!\!\perp A\mid Y$), or sufficiency ($Y\perp\!\!\!\perp A\mid D$). Symbols: #=\#=# = number of, $P=$ probability, $E=$ expected value. Notations: $Y:$ outcome, ${\bf X}:$ features used for model training, $A:$ protected attribute that takes value in the set $\mathcal{A}$, $S:$ model score, $D:$ model classification based on thresholding $S$, ${\bf Z}$: additional set of features.
@@
Independence, sufficiency, and separation provide different perspectives on what it means for a model to be fair. Except under highly restrictive conditions, it is not possible for an algorithm to fulfill all criteria simultaneously [54, 45, 96]. It is therefore critical for healthcare researchers to choose which group fairness considerations are most relevant to the specific context. More specifically, the following pairs of criteria are incompatible in the sense that they cannot generally be simultaneously satisfied: independence and sufficiency, independence and separation, and separation and sufficiency. A basic requirement for any of these pairs to hold is that the outcome and the protected attribute are marginally independent (i.e., $Y\perp\!\!\!\perp A$). In classification problems, this means the probability of being in the positive class is the same across groups - a condition often violated in clinical contexts, where disease prevalence may vary by factors like age or sex. Supplementary Section 3 provides additional mathematical details related to the incompatibilities and also introduces approximate fairness, which allows for small deviations in group fairness metrics to ease these incompatibilities. In Supplementary Section 4, we further explore the practical implications of these incompatibilities through a case study using an openly available EHR dataset.
@@
In contrast to group fairness that targets the average performance of a model across groups, individual fairness ensures that “like cases are treated alike”, an idea grounded in Aristotle’s consistency principles [24, 97]. One of the first works on individual fairness, fairness through awareness (FTA), formalized the notion of similarity by defining metrics that measure the distance between individuals and their predictions. Fairness is achieved if the distance between the prediction for any two individuals is at most the distance between them. A special case of FTA is when the model yields identical decisions for two individuals with all but their protected attributes being the same, a property known as causal discrimination [98].
@@
The implementation of FTA-based approaches critically depends on the choice of distance metrics defined on both predicted outcomes and individuals. For predictions, the choice varies by the type of outcome under study. For example, the Euclidean distance is used for logit transformed predicted probabilities in classification while the absolute difference of the hazard function has been used for survival analysis [99, 100]. However, the metric used to measure the similarity between individuals has been a focus of ongoing research as it encodes exactly how individuals are similar, which relies on an “awareness” of the context [^4] For consistency index proposed by [101], the distance metric for predictions are absolute difference, but it still relies on the choice of metric to define $k$ nearest neighborhood.. In the original FTA paper, the authors suggest leaving the choice of metric to domain experts. As there is currently no agreed upon standard, particularly within health settings, various strategies have been proposed to learn the metric from available data [102, 103, 104]. A scoping review on individual fairness in healthcare provides a comprehensive list of individual fairness methods and corresponding software [31]. The review highlights that individual fairness is just emerging within health-focused applications, likely due to the relative nascency of individual fairness methods developed within the ML literature and the difficult in appropriately quantifying similarity among individuals.
@@
It is important to note that FTA is a counter approach to fairness through unawareness (FTU), which is intended to be a catch all solution to prevent bias by not explicitly including protected attributes into modeling [105] [^5] Note that fairness through awareness implies causal discrimination, but not vice versa.. However, removing protected attributes from a model warrants careful consideration in health-focused applications. First, protected attributes may serve as critical predictors for the outcome of interest, such as age in sepsis or cardiovascular disease prediction and race in cancer screening models [106, 107, 108, 109]. Excluding these predictors can diminish the overall predictive accuracy and impact all individuals adversely or even lead to bias against the majority group. Second, protected attributes are often highly correlated with non-protected attributes. Simply removing protected attributes from the model does not prevent the model from inferring them from other attributes. This is evident in a recent study that found ML models can infer race from medical images [110]. Finally, the quality of the other predictors included in the model may vary across the protected groups (e.g., rates of missing data, measurement error). For instance, in cancer risk prediction, family history serves as an important predictor for disease risk, but was shown to be less reliably documented for Black participants in EHR and therefore less predictive for models developed with these data. [111] found that including race as a predictor significantly improved model performance compared to race-blind algorithms for colorectal cancer. FTU is therefore not always ethical or achievable [112].
@@
Lastly, individual fairness is motivated by the inherent weakness in group fairness criteria that only consider average model performance within groups. That is, there are situations in which group fairness can be satisfied, but individuals within a group can be discriminated against. The incompatibilities among individual and group fairness have been examined from a theoretical perspective in [113, 114, 82, 115]. For example, in predicting ICU admission, if the rate of admission differs significantly across protected attributes, the predictions must be adjusted to achieve statistical parity and obtain an equal rate of positive predictions across the groups. This can cause individuals with similar clinical profiles and medical needs to receive different predictions based solely on their protected attributes, which contradicts the principle of individual fairness. However, individual fairness implies statistical parity when the Wasserstein distance between two groups’ protected attribute distributions is small [24, 116].
@@
We conclude with a brief overview of causal fairness, an increasingly prominent topic in ML that warrants a dedicated review within the health context [117, 118, 119, 120, 121, 31, 75]. As the name suggests, causal fairness focuses on understanding the causal relationships between protected attributes and a model’s decisions [30]. Although causality-based fairness approaches are especially valuable in health applications for mitigating confounding effects and disentangling mechanisms of bias, the use of causal fairness is health is in its early stages.
@@
We introduce one way of incorporating causality into fairness through counterfactuals, which are simply “what-if” statements [122, 123]. In fairness context, the counterfactual is often the unobserved model’s decision that would have happened if the protected attribute had been different [105, 118, 124]. One of the first counterfactual-based fairness criteria, simply called (individual) counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables of interest are held constant [105] [^6] Counterfactual fairness is also viewed as an individual fairness criteria as it is grounded in achieving similar treatment for similar individuals[105].. To better explain this concept, consider an over-simplified view of the world in predicting 30-day hospital readmission. The directed acyclic graph (DAG) in Figure 3 illustrates that the model’s decision for 30-day hospital readmission is directly related to previous diagnoses, age, and prescriptions, while race indirectly influences the decision through its impact on access to healthcare, which in turn influences historical diagnoses [125]. When predicting 30-day hospital readmission, counterfactual fairness is satisfied if a Black patient receives the same prediction they would have received if they were counterfactually of a different race, given all other variables are held constant. The ceteris paribus condition here is crucial, as in reality healthcare utilization may differ across racial communities, which could influence the rate of 30-day hospital readmission in practice. A challenge to causal fairness criteria is that they require a working causal model, the assumptions of which cannot be verified from observational data [105, 126, 118].
@@
Other quantities from the causal inference literature have since been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the model’s decision [127, 40, 128]. Conditional counterfactual fairness is a weaker criteria that considers the total effect of the protected attribute on the decision conditional on a set of variables of interest [105]. In our previous example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the 30-day hospital readmission does not change if their race were counterfactually altered. Another criteria used in health-focused applications is principle fairness, which is a population-level fairness criteria that evaluates decisions among individuals who would be similarly impacted by a model’s decision [129]. Formally, a model satisfies principle fairness if the model’s decision is independent of the protected attribute in principle stratum, which are subgroups with the same values of the counterfactual decision under all possible decisions [129]. More concretely, in our example, principle fairness is achieved if Black and white people have an equal chance of being flagged for 30-day readmission if they have the same likelihood of 30-day readmission (i.e., they belong to the same principal strata). These definitions, along with extensions of the previously introduced group fairness measures — such as counterfactual equalized odds, counterfactual predictive parity, and conditional principal fairness — are formally introduced in Supplementary Section 6. Further work on causal fairness, such as path-specific fairness, direct and indirect discrimination, and interventional fairness, are detailed in [130, 131, 132, 133, 134, 135, 136, 126, 137].
@@
This review explores why models are unfair and various ways fairness has been defined in ML for health. Despite substantial progress in recognizing the importance of bias and in applying group, individual, and causal-based criteria in health applications, there is no consensus on how to appropriately quantify fairness [138, 139]. Each framework comes with unique challenges. Group fairness criteria are easy to implement and interpret, but limited by potential individual fairness violations and by incompatibilities that exist among various criteria. Likewise, individual fairness criteria prioritize similar outcomes for similar individuals to enhance equity in decision making, but can mask disparities at the group level and leave certain populations underserved. Causal fairness, which is gaining traction in health, enables researchers to investigate root causes of unfairness to inform fair decision-making. However, specifying reliable causal models can be challenging in health contexts. Balancing these different frameworks involves weighing the benefits of group-level equity, individual needs, and causal insights [140]. Researchers must remain informed about the full range of available methods and work within interdisciplinary teams (e.g., data scientists, clinicians, ethicists) to make appropriate decisions for their specific application.
It is critical to note that a quantitative understanding of fairness is just one component of operationalizing fairness in ML for health. Establishing common principles, guidelines, and standards that emphasize fairness across the entire model development process is essential, as simply assessing fairness after model training is insufficient [141]. Moreover, most existing fairness definitions were developed outside the health context and do not account for underlying inequities or potential harms to underserved populations [47]. There is an urgent need for context-aware fairness criteria tailored to the unique challenges of health applications [142].
@@
Additionally, health data present inherent challenges, making it difficult - or sometimes impossible - to measure fairness accurately [143, 144, 145, 146, 147, 148]. Various protected attributes (e.g., attributes such as religion, gender identity, and sexual orientation that are missing from Table 2) are often unavailable or inaccurately measured in health datasets due to socio-cultural challenges and lack of standardized data collection practices [149, 62, 150]. While by no means a remedy for inadequate data, there has been recent progress in fairness approaches that operate with limited data on protected attributes that may prove useful in health applications [151, 152]. To date, however, many protected attributes have been omitted from fairness considerations, which underscores the necessity of future methodological work to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [153]. We anticipate future work in these aforementioned directions, particularly in strategies focused on data collection and promoting data equity in health contexts, to establish a strong foundation for fair decision-making.

### v4

@@ Abstract
In predicting hospital readmission, the probability of being readmitted given the model makes that decision and the probability of not being readmitted given the model makes that decision is the same for Black and white patients [raza2023auditing].
Predictive Parity $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospital readmission, the probability of being readmitted given the model makes that decision is the same for Black and white patients [raza2023auditing].
Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all values of the score for males and females [roosli_2022_peeking].
Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all values of the score [roosli_2022_peeking].
@@
Brier Score Parity $\mathbb{E}[(Y-S)^{2}\mid A=a_{0}]=\mathbb{E}[(Y-S)^{2}\mid A=a_{1}],\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to enroll patients into high-risk care management programs, the mean squared error between the score and the label is the same for Black and White patients [obermeyer_dissecting_2019].
Overall Accuracy Equality $P(D=Y\mid A=a_{0})=P(D=Y\mid A=a_{1})\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the probability of a correct classification (correctly classifying a patient as having or not having lung cancer) is the same for males and females [ueda_artificial_2021].
Treatment Equality #$\frac{\#\{D=0,Y=1,A=a_{0}\}}{\#\{D=1,Y=0,A=a_{0}\}}=\frac{\#\{D=0,Y=1,A=a_{1}\}}{\#\{D=1,Y=0,A=a_{1}\}}\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the ratio of the false negatives and false positives is the same for both males and females [ueda_artificial_2021].
Table 3: Common group fairness criteria. Mathematical definitions of group fairness criteria, their interpretation in the context of an example from the literature, and an indication of whether the criterion is a relaxation of independence ($D\perp\!\!\!\perp A$), separation ($D\perp\!\!\!\perp A\mid Y$), or sufficiency ($Y\perp\!\!\!\perp A\mid D$). Symbols: #=#absent\#=# = number of, $P=$ probability, $E=$ expected value. Notations: $Y:$ outcome, ${\bf X}:$ features used for model training, $A:$ protected attribute that takes value in the set $\mathcal{A}$, $S:$ model score, $D:$ model classification based on thresholding $S$, ${\bf Z}$: additional set of features.
@@
Independence, sufficiency, and separation provide different perspectives on what it means for a model to be fair. Except under highly restrictive conditions, it is not possible for an algorithm to fulfill all criteria simultaneously [hardt_equality_2016, chouldechova_fair_2017, corbett-davies_measure_2023]. It is therefore critical for healthcare researchers to choose which group fairness considerations are most relevant to their context. More specifically, the following pairs of criteria are incompatible in the sense that they cannot generally be simultaneously satisfied: independence and sufficiency, independence and separation, and separation and sufficiency. A basic requirement for any of these pairs to hold is that the outcome and the protected attribute are marginally independent (i.e., $Y\perp\!\!\!\perp A$). In classification problems, this means the probability of being in the positive class is the same across groups - a condition often violated in clinical contexts, where disease prevalence may vary by factors like age or sex. Supplementary Section 3 provides additional mathematical details related to the incompatibilities and also introduces approximate fairness, which allows for small deviations in group fairness metrics to relax these incompatibilities. In Supplementary Section 4, we also explore the practical implications of these incompatibilities through a case study using an openly available EHR dataset.
@@
In contrast to group fairness that targets the average performance of a model across groups, individual fairness ensures that “like cases are treated alike”, an idea grounded in Aristotle’s consistency principles [dwork_fairness_2012, fleisher_whats_2021]. One of the first works on individual fairness, fairness through awareness (FTA), formalized the notion of similarity by defining metrics that measure the distance between individuals and their predictions. Fairness is achieved if the distance between the prediction for any two individuals is at most the distance between them. A special case of FTA is when the model yields identical decisions for two individuals with all but their protected attributes being the same, a property known as causal discrimination [galhotra_fairness_2017].
@@
The implementation of FTA-based approaches critically depends on the metrics used to define similarity between individuals and their predictions. For predictions, the choice of metric depends on the type of outcome. For example, the Euclidean distance has been used for logit transformed predicted probabilities in classification while the absolute difference of the hazard function has been used for survival analysis [keya2021equitable, rahman2022fair]. However, the metric used to measure the similarity between individuals has been a focus of ongoing research as it determines how individuals are comparable, which relies on an “awareness” of the context [^4] For consistency index proposed by [zemel_learning_2013], the distance metric for predictions are absolute difference, but it still relies on the choice of metric to define $k$ nearest neighborhood.. In the original FTA paper, the authors suggest leaving the choice of metric to domain experts. As there is currently no agreed upon standard, particularly within health settings, various strategies have been proposed to learn the metric from available data [kim2018fairness, ilvento2019metric, mukherjee2020two]. A scoping review on individual fairness in healthcare provides a comprehensive list of individual fairness methods and corresponding software [anderson2024algorithmic]. The review highlights that individual fairness is just emerging within health-focused applications, likely due to the relative nascency of individual fairness methods developed within the ML literature and the difficult in appropriately quantifying similarity across individuals.
@@
It is important to note that FTA is a counter approach to fairness through unawareness (FTU), which is intended to be a catch all solution to prevent bias by not explicitly including protected attributes into modeling [kusner_counterfactual_2017] [^5] Note that fairness through awareness implies causal discrimination, but not vice versa.. However, removing protected attributes from a model warrants careful consideration in health-focused applications. First, protected attributes may serve as critical predictors for the outcome of interest, such as age in sepsis or cardiovascular disease prediction and race in cancer screening models [damen2016prediction, kim_sex_2018, chicco2020survival, cirillo2020sex]. Excluding these predictors can diminish the overall predictive accuracy and impact all individuals adversely or even lead to bias against the majority group. Second, protected attributes are often highly correlated with non-protected attributes. Simply removing protected attributes from the model does not prevent the model from inferring them from other attributes. This is evident in a recent study that found ML models can infer race from medical images [gichoya_ai_2022]. Finally, the quality of the other predictors included in the model may vary across the protected groups (e.g., rates of missing data, measurement error). For instance, in cancer risk prediction, family history serves as an important predictor for disease risk, but was shown to be less reliably documented for Black participants in self-reported family history data and therefore less predictive for models developed with these data [kupfer_racial_2006, chavez-yenter_association_2022, zink_race_2024]. [zink_race_2024] found that including race as a predictor significantly improved model performance compared to race-blind algorithms for colorectal cancer. FTU is therefore not always ethical, achievable, or desirable [banja2023fairness].
@@
Lastly, individual fairness is motivated by the inherent weakness in group fairness criteria that only consider average model performance within groups. That is, there are situations in which group fairness can be satisfied, but individuals within a group can be discriminated against. The incompatibilities among individual and group fairness have been examined from a theoretical perspective in [friedler_impossibility_2016, binns_apparent_2020, castelnovo_clarification_2022, xu_compatibility_2024]. For example, in predicting ICU admission, if the rate of admission differs significantly across protected attributes, the predictions must be adjusted to achieve statistical parity and obtain an equal rate of positive classifications across the groups. This can cause individuals with similar clinical profiles and medical needs to receive different predictions based solely on their protected attributes, which contradicts the principle of individual fairness. However, individual fairness implies statistical parity when the Wasserstein distance between two groups’ protected attribute distributions is small [dwork_fairness_2012, zhou_group_2022].
@@
We conclude with a brief overview of causal fairness, an increasingly prominent topic in ML that warrants a dedicated review within the health context [loftus2018causal, fawkes2022selection, makhlouf2022survey, carey2022causal, nilforoshan_causal_2022, anderson2024algorithmic, zhang_causal_2024]. As the name suggests, causal fairness focuses on understanding the causal relationships between protected attributes and a model’s decisions [plecko_causal_2022]. Although causality-based fairness approaches are especially valuable in health applications for mitigating confounding effects and disentangling mechanisms of bias, the use of causal fairness is health is in its early stages.
@@
We introduce one way of incorporating causality into fairness through counterfactuals, which are simply “what-if” statements . In fairness context, the counterfactual is often the unobserved model’s decision that would have happened if the protected attribute had been different [kusner_counterfactual_2017, fawkes2022selection, bynum2024new]. One of the first counterfactual-based fairness criteria, simply called (individual) counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables of interest are held constant [kusner_counterfactual_2017] [^6] Counterfactual fairness is also viewed as an individual fairness criteria as it is grounded in achieving similar treatment for similar individuals[kusner_counterfactual_2017].. To better explain this concept, consider an over-simplified view of the world in predicting 30-day hospital readmission. The directed acyclic graph (DAG) in Figure 3 illustrates that the model’s decision for 30-day hospital readmission is directly related to previous diagnoses, age, and prescriptions, while race indirectly influences the decision through its impact on access to healthcare, which in turn influences historical diagnoses [pfohl_counterfactual_2019]. When predicting 30-day hospital readmission, counterfactual fairness is satisfied if a Black patient receives the same prediction they would have received if they were counterfactually of a different race, given all other variables are held constant. The ceteris paribus condition here is crucial, as in reality healthcare utilization may differ within different communities defined by race, which could influence the rate of 30-day hospital readmission in practice. A challenge to causal fairness criteria is that they require a working causal model, the assumptions of which cannot be verified from observational data [kusner_counterfactual_2017, salimi2019interventional, fawkes2022selection].
@@
Numerous other estimands from the causal inference literature have been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the model’s decision [pearl2009causality, mitchell_algorithmic_2021, alves_survey_2023]. Conditional counterfactual fairness is a weaker criteria that considers the total effect of the protected attribute on the decision conditional on a set of variables of interest [kusner_counterfactual_2017]. In our previous example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the 30-day hospital readmission does not change if their race were counterfactually altered. Another criteria used in health-focused applications is principle fairness, which is a population-level fairness criteria that evaluates decisions among individuals who would be similarly impacted the decisions [imai_principal_2023]. Formally, a model satisfies principle fairness if the model’s decision is independent of the protected attribute in principle stratum, which are subgroups with the same values of the counterfactual decision under all possible decisions [imai_principal_2023]. More concretely, in our example, principle fairness is achieved if Black and white people have an equal chance of being flagged for 30-day readmission if they have a similar likelihood of 30-day readmission (i.e., they belong to the same principal strata). These definitions, along with extensions of the previously introduced group fairness measures — such as counterfactual equalized odds, counterfactual predictive parity, and conditional principal fairness — are mathematically defined in Supplementary Section 6. Further work on causal fairness, such as path-specific fairness, direct and indirect discrimination, and interventional fairness, are detailed in [chiappa_path-specific_2019, nabi_fair_2018, wu_pc-fairness_2019, zhang_causal_2017, zhang_mitigating_2018, ogburn_comment_2019, wang_blessings_2019, salimi2019interventional, zuo_interventional_2024].
@@
This review explores why models are unfair and various ways fairness has been defined in ML for health. Despite substantial progress in recognizing the importance of bias and in applying group, individual, and causal-based criteria in health applications, there is no consensus on how to appropriately quantify fairness [zong_medfair_2023, gichoya_equity_2021]. Each framework comes with unique challenges. Group fairness criteria are easy to implement and interpret, but limited by potential individual fairness violations and by incompatibilities that exist among various criteria. Likewise, individual fairness criteria prioritize similar outcomes for similar individuals to enhance equity in decision making, but can mask disparities at the group level and leave certain populations underserved. Causal fairness, which is gaining traction in health, enables researchers to investigate root causes of unfairness to inform fair decision-making. However, specifying reliable causal models can be challenging in health contexts. Balancing these different frameworks involves weighing the benefits of group-level equity, individual needs, and causal insights [schrouff_algorithmic_2022]. Researchers must remain informed about the full range of available methods and work within interdisciplinary teams (e.g., data scientists, clinicians, ethicists) to make appropriate decisions for their specific application. A quantitative understanding of fairness is just one component of operationalizing fairness in ML for health. Establishing common principles, guidelines, and standards that emphasize fairness throughout model development is essential, as simply assessing fairness after model training is insufficient [mccradden_whats_2023].
@@
Health data present inherent and unique challenges, making it difficult - or sometimes impossible - to measure fairness accurately [wick2019unlocking, ji2020can, choi2020fair, zhang2021fairness]. Most existing fairness definitions were developed outside the health context and do not account for underlying inequities or potential harms to underserved populations [mccoy_ensuring_2020]. There is a need for context-aware fairness criteria tailored for health applications [lalor_should_2024]. Additionally, various protected attributes (e.g., attributes such as religion, gender identity, and sexual orientation that are missing from Table 2) are often unavailable or inaccurately measured in health datasets due to sociocultural issues and lack of standardized data collection practices [kronk_transgender_2022, ruberg2020data]. While by no means a remedy for inadequate data, there has been recent progress in fairness approaches that operate with limited data on protected attributes that may prove useful in health applications [zhang_fairlisa_2023, roy_fairness_2024]. To date, however, many protected attributes have been omitted from fairness considerations, which underscores the necessity of future methodological work to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [tomasev2021fairness]. We anticipate future work in these aforementioned directions, particularly in strategies focused on data collection and promoting data equity in health contexts, to establish a strong foundation for fair decision-making.